Widgets Inc common shares are currently trading for $68.25 and the company paid its annual dividend of $1.10 per share. If your required rate of return is 8%, what is the implied growth rate in dividends? (Assume that dividends are expected to grow at a constant rate in perpetuity)
Answer #1

r = [D0 x (1 + g) / P0] + g

0.08 = [$1.10 x (1 + g) / $68.25] + g

0.08 = [0.0161 x (1 + g)] + g

0.08 = 0.0161 + 0.0161g + g

0.0639 = 1.0161g

g = 0.0639 / 1.0161 = 0.0629, or 6.29%